OpenStudy (anonymous):

Help please! Jen wants to save money for the future. Jen invests $700 in an account that pays an interest rate of 6.75% How many years will it take for the account to reach $16,300. Round your answer to the nearest hundredth

OpenStudy (anonymous):

@timbo5lice

OpenStudy (anonymous):

@lochana

OpenStudy (timbo5lice):

ok hold up i gotcha i just need to finish up

OpenStudy (anonymous):

okay(: @timbo5lice

OpenStudy (timbo5lice):

http://cs.selu.edu/~rbyrd/math/continuous/

OpenStudy (lochana):

there is a pattern to tackle 1 year you will have \[700* \frac{106.75}{100} \] 2 year you will have \[(700* \frac{106.75}{100}) \frac{106.75}{100}\] 3 year you will have \[700* (\frac{106.75}{100})^3 \] see the pattern?

OpenStudy (anonymous):

yes kind of

OpenStudy (lochana):

now \[700* (\frac{106.75}{100})^n = 16,300\] \[(\frac{106.75}{100})^n = \frac{16,300}{700}\]

OpenStudy (anonymous):

so the first equation on the last answer u put is the answer so would it be one or two years?? @lochana

OpenStudy (lochana):

I don't remember this question now. was my answer correct?

OpenStudy (lochana):

you will get n = 48.19

OpenStudy (anonymous):

okay so the answer is 48.19? @lochana

OpenStudy (lochana):

I calculated it. I got that

OpenStudy (anonymous):

okay thank you @lochana

OpenStudy (lochana):

you are welcome